graph system of equations and describe it as consistent and independent, consistent or dependent, or inconsistent.
y=-x-2
x+y=-4

Answer by mananth(16949)

y=-x-2
x+y=-4

equation1
when x= 0 y= -2 ( 0 , -2 )
when y = 0 x= -2 ( -2 , 0 )
Equation 2
when x= 0 y= -4 ( 0 , -4 )
when y = 0 x= -4 ( -4 , 0 )
plot the points and draw the line graphs

inconsistent and no solution
m.ananth@hotmail.ca